# LLM Post-Training: A Deep Dive into Reasoning Large Language Models

### v1

@@ 1 Introduction
Contemporary Large Language Models (LLMs) exhibit remarkable capabilities across a vast spectrum of tasks, encompassing not only text generation [1, 2, 3] and question-answering [4, 5, 6, 7], but also sophisticated multi-step reasoning [8, 9, 10, 11]. They power applications in natural language understanding [12, 13, 14, 15, 16, 17], content generation [18, 19, 20, 21, 22, 23, 24, 25], automated reasoning [26, 27, 28, 29], and multimodal interactions [30, 31, 32, 33]. By leveraging vast self-supervised training corpora, these models often approximate human-like cognition [34, 35, 36, 37, 38], demonstrating impressive adaptability in real-world settings.
@@
Adapting LLMs to domain-specific tasks often involves techniques like fine-tuning [54, 55, 56], which enables task-specific learning but risks overfitting and incurs high computational costs. To address these challenges, approaches such as Reinforcement Learning (RL) [57, 58, 59] enhance adaptability by leveraging dynamic feedback and optimizing sequential decision-making. Additionally, advances in scaling techniques, including Low-Rank Adaptation (LoRA[60], adapters, and Retrieval-Augmented Generation (RAG[61, 62, 63], improve both computational efficiency and factual accuracy. These strategies, coupled with distributed training frameworks, facilitate large-scale deployment and further boost the usability of LLMs across diverse applications (Figure 1). Through these targeted post-training interventions, LLMs become better aligned with human intent and ethical requirements, ultimately enhancing their real-world applicability. Below, we summarize key post-training stages.
@@
c) Scaling in LLMs: Scaling is crucial for enhancing the performance and efficiency of LLMs. It helps improve generalization across tasks but introduces significant computational challenges [82, 83]. Balancing performance and resource efficiency requires targeted strategies at inference. Techniques like CoT [8] reasoning and Tree-of-Thought (ToT[84] frameworks enhance multi-step reasoning by breaking down complex problems into sequential or tree-structured steps. Additionally, search-based techniques[85, 86, 87, 88] enable iterative exploration of possible outputs, helping refine responses and ensure higher factual accuracy. These approaches, combined with methods like LoRA [60], adapters, and RAG [61, 62, 89], optimize the model’s ability to handle complex, domain-specific tasks at scale. RAG enhances factual accuracy by dynamically retrieving external knowledge, mitigating limitations of static training data [62, 24, 90]. Distributed training frameworks leverage parallel processing to manage the high computational demands of large-scale models. Test-time scaling optimizes inference by adjusting parameters dynamically based on task complexity [83, 91]. Modifying depth, width, or active layers balances computational efficiency and output quality, making it valuable in resource-limited or variable conditions. Despite advancements, scaling presents challenges such as diminishing returns, longer inference times, and environmental impact, especially when search techniques are performed at test time rather than during training [82]. Ensuring accessibility and feasibility is essential to maintain high-quality, efficient LLM deployment.
@@ 1.1 Prior Surveys
Recent surveys on RL and LLMs provide valuable insights but often focus on specific aspects, leaving key post-training components underexplored [51, 92, 93, 94]. Many works examine RL techniques like Reinforcement Learning from Human Feedback (RLHF[58], Reinforcement Learning from AI Feedback (RLAIF[95], and Direct Preference Optimization (DPO[57], yet they overlook fine-tuning, scaling, and critical benchmarks essential for real-world applications. Furthermore, these studies has not explored the potential of RL even without human annotation supervised finetuning in various frameworks such as DeepSeek R1 with GRPO [59]. Other surveys explore LLMs in traditional RL tasks, such as multi-task learning and decision-making, but they primarily classify LLM functionalities rather than addressing test-time scaling and integrated post-training strategies [96, 97]. Similarly, studies on LLM reasoning [98, 99, 100, 55, 101, 102, 103, 104] discuss learning-to-reason techniques but lack structured guidance on combining fine-tuning, RL, and scaling. The absence of tutorials, along with reviews of software libraries and implementation tools, further limits their practicality. In contrast, this survey offers a comprehensive view of LLM post-training as shown in Figure 1 by systematically covering fine-tuning, RL, and scaling as interconnected optimization strategies. We offer practical resources—benchmarks, datasets, and tutorials—to aid LLM refinement for real-world applications.
@@ 2 Background
$\mathcal{L}_{\text{MLE}}=\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},X).$
Here, $X$ represents the input, such as a prompt or context. $Y=(y_{1},y_{2},...,y_{T})$ is the corresponding target sequence, and $P_{\theta}(y_{t}\mid y_{<t},X)$ denotes the model’s predicted probability for token $y_{t}$, given preceding tokens.
@@
The chain-of-thought reasoning observed in modern LLMs is naturally framed as an RL problem. In this perspective, each intermediate reasoning step is treated as an action contributing to a final answer. The policy gradient update is given by:
@@
Here, we briefly overview pioneering methods that laid the groundwork for applying RL to language generation tasks. These initial efforts train a decision-making model (called a “policy”) by directly adjusting its parameters to maximize rewards. Some policy gradient approaches are explained below:
@@
[Uncaptioned image] Optimizing actions based on long-term rewards, rather than immediate outcomes, remains fundamental in recent LLMs, enabling exploration of multiple reasoning paths.
@@
TABLE I: An overview of reinforcement learning enhanced LLMs, where the notation ‘141B-A39B’ represents a Mixture of Experts (MoE) architecture with a total of 141 billion parameters, out of which 39 billion are actively used during inference.
@@
Advantage Actor-Critic (A2C/A3C). RL methods like REINFORCE [114] rely solely on policy gradients, which suffer from high variance, leading to unstable and inefficient learning. Since the reward signal fluctuates across different trajectories, updates may be noisy, causing slow or erratic convergence. To mitigate this, Actor-Critic methods [152, 153, 154, 155] combine two components: an actor and a critic. The actor is a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that selects actions $a_{t}$ at state $s_{t}$, while the critic is a value function $V_{\phi}(s_{t})$ that evaluates the expected return of a state. The critic provides a more stable learning signal, reducing variance in policy updates and enabling efficient learning in continuous action spaces. Actor updates are guided by the policy gradient theorem, where the advantage function $A(s_{t},a_{t})$ defined in Sec. 2.1, determines how much better an action $a_{t}$ is compared to the expected value of state $s_{t}$. The policy with the learning rate $\alpha$ is updated as:
@@ 3.1 Reward modeling
In practice, a large number of queries $x$ are sampled from real or simulated user requests. Candidate responses $\{y_{j}\}_{j=1}^{m_{x}}$ are generated by either sampling from a base language model or using beam search or other decoding strategies. Human annotators then provide pairwise or ranking feedback on which responses are better (or worse) according to predefined criteria (e.g., quality, correctness, helpfulness, etc). We train a parametric model $R_{\theta}(x,y)$, referred to as the reward model, to map each (query, response) pair $(x,y)$ to a scalar score. The goal is for $R_{\theta}$ to reflect the alignment or preference level, such that:
@@ 3.1.2 Implicit Reward Modeling
Reward Function. Defining a reward function for text generation tasks is an ill-posed problem [166, 167]. The existing RL methods in LLMs either focus on the generation process (Process Reward Modeling) or the outcome (Outcome Reward Modeling), to shape LLM behaviors. We explain these two reward modeling paradigms below.
@@ 3.1.4 Process Reward Modeling
Assigns feedback at intermediate reasoning steps, incentivizing coherent, logically consistent, and well-structured chains of thought. This approach is particularly valuable for tasks involving mathematical derivations, legal arguments, or code debugging, in which the path to the answer is as significant as the final statement. In such problems, the reward assigned at individual steps encourages transparency and robust step-by-step reasoning. However, it requires a more complex annotation process, e.g., requires “gold” reasoning steps or partial credit scoring. Process rewards can be combined with outcome rewards for a strong multi-phase training signal.
[Uncaptioned image] PRM with last-step aggregation beats ORM.
@@ 3.2.1 Odds Ratio Preference Optimization (ORPO)
A simplest method ORPO [168] which directly optimizing a policy from pairwise human preferences. Instead of first learning a separate reward model and then running standard RL, ORPO updates the policy to increase the likelihood of preferred responses (according to human labels) relative to dispreferred ones. The key idea is to look at the odds ratio:
@@ 3.2.2 Proximal Policy Optimization (PPO) in LLMs
[Uncaptioned image] The KL penalty ensure policy retains the base model’s linguistic coherence and avoids degenerate outputs.
@@ 3.2.6 Direct Preference Optimization (DPO)
Perplexity Filtering for On-Distribution Data. To ensure DPO training data is on-distribution (aligned with $\rho$), responses are filtered using perplexity. The perplexity of a response $y=(y_{1},y_{2},\dots,y_{T})$ is defined as:
@@ 3.3.1 Cold-Start RL Phase
[Uncaptioned image] CoT examples before RL training provide a stronger foundation for reasoning tasks, leading to more robust and interpretable outputs.
@@ 5. Integrated Development Ecosystems
TABLE II: Comprehensive Overview of Modern LLM Methods and Frameworks.
@@ 5.1 Beam Search
While modern LLMs often favor stochastic sampling (e.g., temperature sampling) to promote diversity in generated text, beam search is still a valuable technique for structured reasoning problems. For example, the Tree-of-Thoughts framework [84] allows plugging in different search algorithms to explore a tree of possible ‘thoughts’ or reasoning steps; one instantiation uses a beam search (with beam width $b$) to maintain the $b$ most promising states at each reasoning step. Here, beam search is used to systematically explore solution steps for tasks like mathematical puzzles and planning problems, pruning less promising reasoning branches and thus improving the model’s problem-solving accuracy. Beam search remains a strong baseline for test-time reasoning when one wants the model to output the single most likely reasoning path or answer under the model’s learned distribution.
@@ 5.2 Best-of-N Search (Rejection Sampling)
[Uncaptioned image] Using process reward models with beam search vs best-of-N depends on the difficulty and compute budget.
@@ 5.4 Chain-of-thought prompting
[Uncaptioned image] Finetune models to revise answers sequentially, leveraging previous attempts. Sequential revisions excel on easier questions, while parallel sampling (exploration) benefits harder ones
@@ 5.8 Confidence-based Sampling
Confidence-based strategies have been incorporated at inference time e.g., a tree-based search for LLM generation [248] assigns each possible completion (leaf) a confidence score. The algorithm samples leaves in proportion to these confidence scores to decide which paths to extend. Similarly, some reasoning approaches use the model’s estimated likelihood of an answer to decide when to halt or whether to ask a follow-up question – essentially if the model’s confidence is low, it might trigger further reasoning (a form of self-reflection). Confidence-based selection is also used in ensemble settings: e.g., an LLM may generate multiple answers and a secondary model evaluates the confidence of each answer being correct, picking the answer with the highest confidence. This was explored in tasks like medical Q&A, where an LLM gave an answer and a confidence score, and only high confidence answers were trusted or returned [250].
@@ 5.11 Monte Carlo Tree Search
[Uncaptioned image] Test-time compute is not a 1-to-1 replacement for pretraining but offers a viable alternative in many cases.
@@ 5.12 Chain-of-Action-Thought reasoning
[Uncaptioned image] A smaller model with test-time compute can outperform a 14× larger model on easy/intermediate questions when inference tokens (Y) are limited (e.g., self-improvement settings).
@@ 6 Benchmarks for LLM Post-training Evaluation
Datasets Domain Type #Samples Evaluation Criteria Reasoning Benchmarks MATH  [269] Math Reasoning Pointwise 7,500 Step-by-step solutions GSM8K  [270] Math Reasoning Pointwise 8.5K Multi-step reasoning MetaMathQA [271] Math Reasoning Pointwise 40K+ Self-verification, FOBAR WorldTree V2  [272] Science QA Pointwise 1,680 Multi-hop explanations PangeaBench  [273] Multimodal Reasoning Pairwise 47 Langs. Cultural understanding MMMU  [274] Science/Math Pointwise College-Level Physics, Chemistry, Bilingual TruthfulQA  [275] QA/Reasoning Pointwise N/A Truthfulness MathInstruct  [276] Math Reasoning Pointwise 262K Correctness MMLU  [277, 278] Multitask Reasoning Pointwise 57 Tasks Broad knowledge evaluation MMLU-Fairness  [277] Fairness/Reasoning Pointwise N/A Bias/Equity Analysis DROP  [279] Reading/Reasoning Pointwise 96K Discrete reasoning over paragraphs BBH  [175] Hard Reasoning Pairwise N/A Complex logical problem-solving VRC-Bench [187] Multimodal Reasoning Pairwise N/A Visual Reasoning and Classification RL Alignment Benchmarks HelpSteer  [280] RL Alignment Pairwise 37K+ Multi-attribute scoring Anthropic HH-RLHF  [121] RL Alignment Pairwise 42.5K Harmlessness alignment UltraFeedback  [281] RL Alignment Pairwise 64K Instruction-following, Truthfulness D4RL  [282] RL/Control Pointwise N/A Offline RL across domains Meta-World  [283] RL/Control Pointwise N/A Multi-task robotic RL MineRL  [284] RL/Games Pairwise N/A Imitation learning, rewards Multilingual Evaluation CulturaX  [285] Multilingual Pointwise 6.3T Deduplication, Quality PangeaIns  [286] Multilingual Pointwise 6M Multilingual instructions TydiQA [287] Multilingual Pointwise N/A Cross-lingual QA XGLUE  [288] Multilingual Pointwise N/A Cross-lingual language tasks MM-Eval [289] Multilingual Pairwise 4,981 Task-oriented multilingual QA ALM-Bench [289] Multilingual QA Pointwise N/A Multilingual Evaluation General Comprehension Benchmarks BigBench [290] General Comprehension Pointwise 200+ Tasks Broad multi-domain evaluation Chatbot Arena [291] Comprehension Pairwise 33K User preference MTBench  [291] Comprehension Pairwise 3K Multi-turn conversations RewardBench [167] Comprehension Pairwise 2,998 User preference General Comprehension Benchmarks ConvAI2  [292] Dialogue Pointwise N/A Engagingness, Consistency MultiWOZ [293] Dialogue Pointwise N/A Task success, Coherence Trec DL21&22 [294, 295] Search Pointwise 1,549/2,673 Relevance scoring BEIR [296] Search Pointwise 18 Datasets Information retrieval Story & Recommendation Benchmarks HANNA  [297] Story Pointwise 1,056 Relevance, Coherence, Complexity StoryER [298] Story Pairwise 100K User preference-based ranking PKU-SafeRLHF [299] Values Pairwise 83.4K Helpfulness, Harmlessness Cvalue [300] Values Pairwise 145K Safety, Responsibility NaturalInst. [301, 302] Instruction Tuning Pointwise 1,600+ Instruction-following evaluation

### v2

@@ 1 Introduction
Contemporary Large Language Models (LLMs) exhibit remarkable capabilities across a vast spectrum of tasks, encompassing not only text generation [1, 2, 3] and question-answering [4, 5, 6, 7], but also sophisticated multi-step reasoning [8, 9, 10, 11]. They power applications in natural language understanding [12, 13, 14, 15, 16, 17], content generation [18, 19, 20, 21, 22, 23, 24, 25], automated reasoning [26, 27, 28, 29], and multimodal interactions [30, 31, ashraf2024fate, 33]. By leveraging vast self-supervised training corpora, these models often approximate human-like cognition [34, 35, 36, 37, 38], demonstrating impressive adaptability in real-world settings.
@@
Adapting LLMs to domain-specific tasks often involves techniques like fine-tuning [54, 55, 56], which enables task-specific learning but risks overfitting and incurs high computational costs. To address these challenges, approaches such as Reinforcement Learning (RL) [57, 58, 59] enhance adaptability by leveraging dynamic feedback and optimizing sequential decision-making. Additionally, advances in scaling techniques, including Low-Rank Adaptation (LoRA[60], adapters [365, tafjord2021proofwritergeneratingimplicationsproofs], and Retrieval-Augmented Generation (RAG[61, 62, 63], improve both computational efficiency and factual accuracy. These strategies, coupled with distributed training frameworks, facilitate large-scale deployment and further boost the usability of LLMs across diverse applications (Figure 1). Through these targeted post-training interventions, LLMs become better aligned with human intent and ethical requirements, ultimately enhancing their real-world applicability. Below, we summarize key post-training stages.
@@
c) Test Time Scaling in LLMs: Test Time Scaling is optimizing model performance and efficiency without altering the core architecture. It enables better generalization while minimizing computational overhead. It is crucial for enhancing the performance and efficiency of LLMs. It helps improve generalization across tasks but introduces significant computational challenges [82, 83]. Balancing performance and resource efficiency requires targeted strategies at inference. Techniques like CoT [8] reasoning and Tree-of-Thought (ToT[84] frameworks enhance multi-step reasoning by breaking down complex problems into sequential or tree-structured steps. Additionally, search-based techniques[85, 86, 87, 88] enable iterative exploration of possible outputs, helping refine responses and ensure higher factual accuracy. These approaches, combined with methods like LoRA [60], adapters, and RAG [61, 62, 89], optimize the model’s ability to handle complex, domain-specific tasks at scale. RAG enhances factual accuracy by dynamically retrieving external knowledge, mitigating limitations of static training data [62, 24, 90]. Distributed training frameworks leverage parallel processing to manage the high computational demands of large-scale models. Test-time scaling optimizes inference by adjusting parameters dynamically based on task complexity [83, 91]. Modifying depth, width, or active layers balances computational efficiency and output quality, making it valuable in resource-limited or variable conditions. Despite advancements, scaling presents challenges such as diminishing returns, longer inference times, and environmental impact, especially when search techniques are performed at test time rather than during training [82]. Ensuring accessibility and feasibility is essential to maintain high-quality, efficient LLM deployment.
@@ 1.1 Prior Surveys
Recent surveys on RL and LLMs provide valuable insights but often focus on specific aspects, leaving key post-training components underexplored [51, 92, 93, 94]. Many works examine RL techniques like Reinforcement Learning from Human Feedback (RLHF[58], Reinforcement Learning from AI Feedback (RLAIF[95], and Direct Preference Optimization (DPO[57], yet they overlook fine-tuning, scaling, and critical benchmarks essential for real-world applications. Furthermore, these studies have not explored the potential of RL even without human annotation supervised finetuning in various frameworks such as DeepSeek R1 with GRPO [59]. Other surveys explore LLMs in traditional RL tasks, such as multi-task learning and decision-making, but they primarily classify LLM functionalities rather than addressing test-time scaling and integrated post-training strategies [96, 97]. Similarly, studies on LLM reasoning [98, 99, 100, 55, 101, 102, 103, 104] discuss learning-to-reason techniques but lack structured guidance on combining fine-tuning, RL, and scaling. The absence of tutorials, along with reviews of software libraries and implementation tools, further limits their practicality. In contrast, this survey offers a comprehensive view of LLM post-training as shown in Figure 1 by systematically covering fine-tuning, RL, and scaling as interconnected optimization strategies. We offer practical resources—benchmarks, datasets, and tutorials—to aid LLM refinement for real-world applications.
@@ 2 Background
$\mathcal{L}_{\text{MLE}}=-\sum_{t=1}^{T}\log P_{\theta}(y_{t}\mid y_{<t},X).$
Here, $X$ represents the input, such as a prompt or context. $Y=(y_{1},y_{2},...,y_{T})$ is the corresponding target output sequence, and $P_{\theta}(y_{t}\mid y_{<t},X)$ denotes the model’s predicted probability for token $y_{t}$, given preceding tokens.
@@
The chain-of-thought reasoning employed in modern LLMs is naturally framed as an RL problem. In this perspective, each intermediate reasoning step is treated as an action contributing to a final answer. The objective function $J(\pi_{\theta})$ represents the expected reward of the policy $\pi_{\theta}$, capturing how well the model performs over multiple reasoning steps. The policy gradient update is given by:
@@
Here, we briefly overview pioneering methods that laid the groundwork for applying RL to language generation tasks. These initial efforts train a decision-making model (policy ($p_{\theta}$)) by directly adjusting its parameters to maximize rewards. Some policy gradient approaches are explained below:
@@
[Uncaptioned image] Optimizing actions based on long-term rewards, which account for the cumulative benefits of a sequence of reasoning steps rather than just immediate outcomes, is fundamental in recent LLMs. This approach allows models to explore multiple reasoning paths more effectively.
@@
TABLE I: An overview of reinforcement learning-enhanced LLMs, where ’141B-A39B’ denotes a Mixture of Experts (MoE) model with 141 billion total parameters, of which 39 billion are utilized during inference. TTS stands for Test-Time Scaling.
@@
Advantage Actor-Critic (A2C/A3C). RL methods like REINFORCE [114] rely solely on policy gradients, which suffer from high variance, leading to unstable and inefficient learning. Since the reward signal fluctuates across different trajectories, updates may be noisy, causing slow or erratic convergence. To mitigate this, Actor-Critic methods [152, 153, 154, 155] combine two components as follows: an actor and a critic. The actor is a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that selects actions $a_{t}$ at state $s_{t}$, while the critic is a value function $V_{\phi}(s_{t})$ that evaluates the expected return of a state. The critic provides a more stable learning signal, reducing variance in policy updates and enabling efficient learning in continuous action spaces. Actor updates are guided by the policy gradient theorem, where the advantage function $A(s_{t},a_{t})$ defined in Sec. 2.1, determines how much better an action $a_{t}$ is compared to the expected value of state $s_{t}$. The policy with the learning rate $\alpha$ is updated as:
@@ 3.1 Reward modeling
In practice, a large number of queries $x$ are sampled from real or simulated user requests. Candidate responses $\{y_{j}\}_{j=1}^{m_{x}}$ are generated by either sampling from a base language model or using beam search or other decoding strategies. Human annotators then provide pairwise or ranking feedback on which responses are better (or worse) according to predefined criteria (e.g., quality, correctness, helpfulness, etc). We train a parametric model Reward Model ($R_{\theta}(x,y)$), referred to as the reward model, to map each (query, response) pair $(x,y)$ to a scalar score. The goal is for $R_{\theta}$ to reflect the alignment or preference level, such that:
@@ 3.1.2 Implicit Reward Modeling
Reward Function. Defining a reward function for text generation tasks is an ill-posed problem [166, 167]. The existing RL methods in LLMs either focus on the generation process outcome (Outcome Reward Modeling) or the (Process Reward Modeling), to shape LLM behaviors. We explain these two reward modeling paradigms below.
@@ 3.1.4 Process Reward Modeling
Assigns feedback at intermediate reasoning steps, incentivizing coherent, logically consistent, and well-structured chains of thought. This approach is particularly valuable for tasks involving mathematical derivations, legal arguments, or code debugging, in which the path to the answer is as significant as the final statement. In such problems, the reward assigned in individual steps encourages transparency and robust step-by-step reasoning. However, it requires a more complex annotation process, e.g., requires “gold” reasoning steps or partial credit scoring. Process rewards can be combined with outcome rewards for a strong multi-phase training signal.
[Uncaptioned image] Policy Reward Modeling (PRM) with last-step aggregation outperforms Outcome Reward Modeling (ORM) by leveraging final-step evaluations to optimize policy updates more effectively.
@@ 3.2.1 Odds Ratio Preference Optimization (ORPO)
The simplest method is ORPO [168] which directly optimizing a policy from pairwise human preferences. Instead of first learning a separate reward model and then running standard RL, ORPO updates the policy to increase the likelihood of preferred responses (according to human labels) relative to dispreferred ones. The key idea is to look at the odds ratio:
@@ 3.2.2 Proximal Policy Optimization (PPO) in LLMs
[Uncaptioned image] The KL penalty is a regularization, which ensure policy retains the base model’s linguistic coherence and avoids degenerate outputs.
@@ 3.2.6 Direct Preference Optimization (DPO)
Perplexity Filtering for Out-of-Distribution Data. To ensure DPO training data is on-distribution (aligned with $\rho$), responses are filtered using perplexity. The perplexity of a response $y=(y_{1},y_{2},\dots,y_{T})$ is defined as:
@@ 3.3.1 Cold-Start RL Phase
[Uncaptioned image] Providing CoT reasoning traces before RL training establishes a stronger foundation for reasoning tasks, enhancing both robustness and interpretability of outputs.
@@ 5. Integrated Development Ecosystems
TABLE II: Comprehensive Overview of Methods and Frameworks employed in Modern LLMs
@@ 5.1 Beam Search
While modern LLMs often favor stochastic sampling (e.g., temperature sampling) to promote diversity in generated text, beam search is still a valuable technique for structured reasoning problems. For example, the Tree-of-Thoughts framework [84] allows plugging in different search algorithms to explore a tree of possible ‘thoughts’ or reasoning steps; usually a beam search (with beam width $b$) is used to maintain the $b$ most promising states at each reasoning step. Here, beam search is used to systematically explore solution steps for tasks like mathematical puzzles and planning problems, pruning less promising reasoning branches and thus improving the model’s problem-solving accuracy. Beam search remains a strong baseline for test-time reasoning when one wants the model to output the single most likely reasoning path or answer under the model’s learned distribution.
@@ 5.2 Best-of-N Search (Rejection Sampling)
[Uncaptioned image] Choice of either process reward models with beam search vs best-of-N depends on the difficulty and compute budget.
@@ 5.4 Chain-of-thought prompting
[Uncaptioned image] Fine-tuning models to revise answers sequentially allows them to build on previous attempts, improving accuracy over time. This approach is particularly effective for easier questions, while parallel sampling (exploration) proves more beneficial for harder ones.
@@ 5.8 Confidence-based Sampling
Confidence-based strategies have been incorporated at inference time e.g., a tree-based search for LLM generation [248] assigns each possible completion (leaf) a confidence score. The algorithm samples leaves in proportion to these confidence scores to decide which paths to extend [272]. Similarly, some reasoning approaches use the model’s estimated likelihood of an answer to decide when to halt or whether to ask a follow-up question – essentially if the model’s confidence is low, it might trigger further reasoning (a form of self-reflection). Confidence-based selection is also used in ensemble settings: e.g., an LLM may generate multiple answers and a secondary model evaluates the confidence of each answer being correct, picking the answer with the highest confidence. This was explored in tasks like medical Q&A, where an LLM gave an answer and a confidence score, and only high confidence answers were trusted or returned [250].
@@ 5.11 Monte Carlo Tree Search
[Uncaptioned image] Test-time compute is not a 1-to-1 replacement for pretraining but, offers a viable alternative in many cases.
@@ 5.12 Chain-of-Action-Thought reasoning
[Uncaptioned image] A smaller model with test-time compute can outperform a 14× larger model on easy/intermediate questions, when inference tokens (Y) are limited (e.g., self-improvement settings).
@@ 6 Benchmarks for LLM Post-training Evaluation
Datasets Domain Type #Samples Evaluation Criteria Reasoning Benchmarks MATH  [269] Math Reasoning Pointwise 7,500 Step-by-step solutions GSM8K  [270] Math Reasoning Pointwise 8.5K Multi-step reasoning MetaMathQA [271] Math Reasoning Pointwise 40K+ Self-verification, FOBAR WorldTree V2  [272] Science QA Pointwise 1,680 Multi-hop explanations PangeaBench  [273] Multimodal Reasoning Pairwise 47 Langs. Cultural understanding MMMU  [274] Science/Math Pointwise College-Level Physics, Chemistry, Bilingual TruthfulQA  [275] QA/Reasoning Pointwise N/A Truthfulness MathInstruct  [276] Math Reasoning Pointwise 262K Correctness MMLU  [277, 278] Multitask Reasoning Pointwise 57 Tasks Broad knowledge evaluation MMLU-Fairness  [277] Fairness/Reasoning Pointwise N/A Bias/Equity Analysis DROP  [279] Reading/Reasoning Pointwise 96K Discrete reasoning over paragraphs BBH  [175] Hard Reasoning Pairwise N/A Complex logical problem-solving VRC-Bench [187] Multimodal Reasoning Pairwise N/A Visual Reasoning and Classification RL Alignment Benchmarks HelpSteer  [280] RL Alignment Pairwise 37K+ Multi-attribute scoring Anthropic HH-RLHF  [121] RL Alignment Pairwise 42.5K Harmlessness alignment UltraFeedback  [281] RL Alignment Pairwise 64K Instruction-following, Truthfulness D4RL  [282] RL/Control Pointwise N/A Offline RL across domains Meta-World  [283] RL/Control Pointwise N/A Multi-task robotic RL MineRL  [284] RL/Games Pairwise N/A Imitation learning, rewards Multilingual Evaluation CulturaX  [285] Multilingual Pointwise 6.3T Deduplication, Quality PangeaIns  [286] Multilingual Pointwise 6M Multilingual instructions TydiQA [287] Multilingual Pointwise N/A Cross-lingual QA XGLUE  [288] Multilingual Pointwise N/A Cross-lingual language tasks MM-Eval [289] Multilingual Pairwise 4,981 Task-oriented multilingual QA ALM-Bench [289] Multilingual QA Pointwise N/A Multilingual Evaluation Dialogue and Search Benchmarks BigBench [290] General Comprehension Pointwise 200+ Tasks Broad multi-domain evaluation Chatbot Arena [291] Comprehension Pairwise 33K User preference MTBench  [291] Comprehension Pairwise 3K Multi-turn conversations RewardBench [167] Comprehension Pairwise 2,998 User preference General Comprehension Benchmarks ConvAI2  [292] Dialogue Pointwise N/A Engagingness, Consistency MultiWOZ [293] Dialogue Pointwise N/A Task success, Coherence Trec DL21&22 [294, 295] Search Pointwise 1,549/2,673 Relevance scoring BEIR [296] Search Pointwise 18 Datasets Information retrieval Story & Recommendation Benchmarks HANNA  [297] Story Pointwise 1,056 Relevance, Coherence, Complexity StoryER [298] Story Pairwise 100K User preference-based ranking PKU-SafeRLHF [299] Values Pairwise 83.4K Helpfulness, Harmlessness Cvalue [300] Values Pairwise 145K Safety, Responsibility NaturalInst. [301, 302] Instruction Tuning Pointwise 1,600+ Instruction-following evaluation